# CITYWALK: Enhancing LLM-Based C++ Unit Test Generation via Project-Dependency Awareness and Language-Specific Knowledge

### v1

@@ 4.1.2. Bad Case Breakdown
We further conduct an in-depth investigation into the common error categories of the failed test cases generated by CITYWALK. Specifically, we manually analyze the output error messages associated with these failed test cases and summarize the corresponding error categories. In total, CITYWALK produces 1329 failed test cases across eight projects. Our analysis reveals the following findings based on the statistical results:
@@
    The two most prevalent categories of errors in the failed test cases are compilation errors and execution errors. As illustrated in Figure 7, we subdivide the two high-level categories into distinct sub-categories. Due to space constraints, we only present the errors that are observed more than ten times. The most common sub-category of compilation errors is Access Error, which occurs when LLMs frequently generate test code that attempts to invalidly access private variables or methods. The most common sub-category of execution errors is Test Setup Error, primarily due to LLMs facing challenges in accurately inferring the mock object configurations, as well as generating valid test data inputs.
@@
    The complexity of C++ language features poses significant challenges for CITYWALK in generating accurate test cases. In the design of CITYWALK, we integrate language-specific domain knowledge derived from empirical observations to address common errors, thereby enhancing the performance of the LLM to some extent. However, since the benchmark is collected from a diverse range of real-world projects, project-specific issues may still impede the generation of correct and high-coverage test code. For instance, many functionalities in the large basic software system leveldb rely heavily on external interactions, such as disk I/O operations and file content checks. This dependence often compels LLMs to redefine existing methods in dependent classes or to invoke undefined methods when constructing I/O streams, leading to errors such as Multiple Definition Error and Undefined Symbol Error, which result in low evaluation metrics. Additionally, leveldb involves advanced features like synchronization locks, which significantly increase the complexity and challenge the test generation process. Moreover, the absence of an automated and strict exception handling mechanism, akin to that in Java, complicates the LLM’s ability to generate correct test code with exception handling logic. The most frequently occurring assertion errors can be traced back to the lack of supplementary post-processing techniques, which could help prompt LLMs to rectify incorrect assertions in the generated test cases. We are concerned that providing LLMs with error messages that include expected outputs may encourage them to replicate those outputs directly, thus manipulating the compiler into successful execution of the generated test cases. We will discuss these false positive cases in detail in Section 5.2.
Figure 8. The Ablation Study for CITYWALK in Terms of the Coverage Metrics.
[…]
@@
Answer to RQ1: In conclusion, CITYWALK exhibits a marked superiority over the LLM-based baselines across all evaluation metrics, highlighting its effectiveness in C++ unit test generation.
@@ 4.2. Answering RQ2
To answer this question, we conduct a series of ablation experiments to assess the impact of different designed components within CITYWALK, including configuration dependencies ${\mathtt{Dep_{c}}}$, cross-file data dependencies $\mathtt{Dep_{d}}$, intention contexts $\mathtt{Context_{intent}}$, guidelines of language-specific domain knowledge $\mathtt{Guideline_{DK}}$, step-by-step instructions $\mathtt{PROMPT_{step}}$, and post-processing techniques $\mathtt{Fix_{rule+prompt}}$. To ensure fairness in comparisons, the parameter configurations align with those described in Section 3.5. Specifically, we perform ablation experiments by removing one component at a time and analyze the performance contribution of each component regarding the coverage metrics $\mathbf{Cov_{L}}$ and $\mathbf{Cov_{B}}$. Figure 8 illustrates the comparison results of CITYWALK against the ablated models across all evaluated projects. When compared to CITYWALK, each ablation model exhibits varying degrees of decline in terms of $\mathbf{Cov_{L}}$ and $\mathbf{Cov_{B}}$, indicating that each designed component contributes to the overall enhancement in generating high-quality unit test cases. Table 8 and Table 9 show the performance contribution results (denoted as the performance degradation values). The greater the coverage scores decline, the larger the contribution of that component. The Avg. $\downarrow$ columns present the average results of all projects. According to the statistical results presented in Table 8 and Table 9, the top three components that contribute the most to CITYWALK are ${\mathtt{Dep_{c}}}$, ${\mathtt{Dep_{d}}}$, and $\mathtt{Context_{intent}}$. As discussed in Section 4.1.1, ${\mathtt{Dep_{c}}}$ plays a crucial role in guiding LLMs to generate syntactically correct test code by directly incorporating relevant configuration dependencies. Therefore, projects such as hjson-cpp, tinyxml2, re2, and json.cpp, which do not use gtest, would benefit from ${\mathtt{Dep_{c}}}$ by helping the LLMs understand the absence of such a testing framework, thereby preventing the generation of non-compilable test cases. Figure 9 illustrates how extracted cross-file data dependencies are used as guidance for LLMs to generate correct test cases. By supplementing the key $\mathtt{Dep_{d}}$ (i.e., the initialization constructor of Json) from the json.h file, LLMs can avoid errors that would otherwise arise from using non-existent functions (i.e., setNumber). Meanwhile, Figure 10 illustrates how $\mathtt{Context_{intent}}$ can guide LLMs in generating correct test cases for the Failed Test Case ➋ within Figure 1. Specifically, the retrieved ParseTag method within the singledocparser class provides an invocation example of the focal method, while the retrieved _Tag method within the emittermanip header file offers an initialization example of the focal method. These examples significantly aid LLMs in understanding the usage patterns of the focal method, contributing to the generation of functionally correct assertions and enhancing code coverage.
@@
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\downarrow$
CITYWALK 27.94% 77.61% 44.46% 42.56% 56.60% 25.25% 50.71% 42.15%
   w/o ${\mathtt{Dep_{c}}}$27.94% -77.31% -12.08% -42.56% -3.82% -1.36% -50.71% -1.81% -27.20%
   w/o ${\mathtt{Dep_{d}}}$0.82% -32.42% -14.50% -7.07% -15.26% -8.69% -19.04% -13.03% -13.85%
   w/o ${\mathtt{Context_{intent}}}$12.36% -29.84% -8.72% -8.21% -11.21% -5.11% -6.59% -9.70% -11.47%
   w/o ${\mathtt{Guideline_{DK}}}$16.74% -7.40% -11.48% -7.87% -7.04% -2.69% -3.59% -5.59% -7.80%
   w/o ${\mathtt{PROMPT_{step}}}$5.46% -12.97% -12.33% -9.86% -8.42% -8.91% -12.48% -7.59% -9.75%
   w/o ${\mathtt{Fix_{rule+prompt}}}$5.24% -12.51% -14.95% -8.48% -10.76% -1.51% -1.27% -4.26% -7.37%
@@
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap Avg. $\downarrow$
CITYWALK 24.29% 65.66% 35.56% 35.64% 49.72% 24.64% 46.61% 37.82%
   w/o ${\mathtt{Dep_{c}}}$24.29% -65.66% -10.04% -35.64% -3.38% -1.50% -46.61% -1.70% -23.60%
   w/o ${\mathtt{Dep_{d}}}$1.99% -30.69% -11.48% -7.94% -14.60% -8.75% -19.33% -13.93% -13.59%
   w/o ${\mathtt{Context_{intent}}}$14.56% -28.48% -7.00% -9.87% -9.60% -5.17% -6.83% -9.59% -11.39%
   w/o ${\mathtt{Guideline_{DK}}}$15.51% -8.07% -8.70% -9.02% -5.83% -2.75% -3.72% -5.70% -7.41%
   w/o ${\mathtt{PROMPT_{step}}}$5.97% -13.60% -10.19% -9.90% -6.16% -8.97% -12.49% -7.59% -9.36%
   w/o ${\mathtt{Fix_{rule+prompt}}}$6.12% -9.68% -8.52% -9.60% -$3$11% -0.64% -3.26% -8.91%

### v2

@@ 4.1.2. Bad Case Breakdown
We further conduct an in-depth investigation into the common error categories of the failed test cases generated by CITYWALK. Specifically, we manually analyze the output error messages associated with these failed test cases and summarize the corresponding error categories. In total, CITYWALK produces 1567 errors across 1021 failed test cases in all ten projects. Since a single test case can contain multiple errors, the total error count exceeds the number of failed test cases. Our analysis reveals the following findings based on the statistical results:
@@
    The two most prevalent categories of errors in the failed test cases are compilation errors and execution errors. As illustrated in Figure 8, we subdivide the two high-level categories into distinct sub-categories. The most common sub-category of compilation errors is Access Error, which occurs when LLMs frequently generate test code that attempts to invalidly access private variables or methods. The most common sub-category of execution errors is Test Setup Error, primarily due to LLMs facing challenges in accurately inferring the mock object configurations, as well as generating valid test data inputs.
@@
    The complexity of C++ language features poses significant challenges for CITYWALK in generating accurate test cases. In the design of CITYWALK, we integrate language-specific domain knowledge derived from empirical observations to address common errors, thereby enhancing the performance of the LLM to some extent. However, since the benchmark is collected from a diverse range of real-world projects, project-specific issues may still impede the generation of correct and high-coverage test code. For instance, many functionalities in the large basic software system leveldb rely heavily on external interactions, such as disk I/O operations and file content checks. This dependence often compels LLMs to redefine existing methods in dependent classes or to invoke undefined methods when constructing I/O streams, leading to errors such as Multiple Definition Error and Undefined Symbol Error, which result in low evaluation metrics. Additionally, leveldb involves advanced features like synchronization locks, which significantly increase the complexity and challenge the test generation process. Moreover, the absence of an automated and strict exception handling mechanism, akin to that in Java, complicates the LLM’s ability to generate correct test code with exception handling logic. The most frequently occurring assertion errors can be traced back to the lack of supplementary post-processing techniques (e.g., post-hoc fix), which could help prompt LLMs to rectify incorrect assertions in the generated test cases. We are concerned that providing LLMs with error messages that include expected outputs may encourage them to replicate those outputs directly, thus manipulating the compiler into successful execution of the generated test cases. We will discuss these false positive cases in detail in Section 5.4.
Table 10. Generalization Results of CITYWALK on Different LLMs in Terms of Compilation Success Rate (CSR)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg. $\uparrow$
CodeGeeX4 0.00% 6.95% 38.98% 0.09% 4.22% 11.75% 10.70% 35.75% 50.00% 51.35% 19.76%
CodeGeeX4 w/ CITYWALK 43.59% 42.11% 54.64% 4.30% 31.45% 29.12% 36.45% 42.44% 61.11% 62.22%
DeepSeek-V3 0.00% 0.00% 17.79% 0.00% 7.35% 13.55% 25.14% 51.93% 18.75% 68.34% 37.94%
DeepSeek-V3 w/ CITYWALK 74.62% 72.67% 43.45% 47.83% 41.56% 42.53% 53.89% 62.34% 70.11% 73.24%
GPT-3.5 0.00% 0.00% 19.34% 5.45% 5.00% 9.64% 31.16% 28.57% 52.22% 72.80% 23.66%
GPT-3.5 w/ CITYWALK 31.78% 72.96% 33.69% 20.05% 27.53% 32.45% 45.23% 37.34% 86.67% 73.07%
Table 11. Generalization Results of CITYWALK on Different LLMs in Terms of Execution Pass Rate (EPR)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg. $\uparrow$
CodeGeeX4 0.00% 6.00% 2.76% 0.09% 3.82% 0.36% 8.96% 15.54% 40.54% 41.89% 21.36%
CodeGeeX4 w/ CITYWALK 30.77% 37.92% 42.92% 2.15% 26.78% 25.56% 21.78% 32.34% 55.56% 57.78%
DeepSeek-V3 0.00% 0.00% 13.78% 0.00% 6.69% 12.02% 24.02% 39.91% 18.75% 66.29% 33.43%
DeepSeek-V3 w/ CITYWALK 67.69% 64.92% 31.82% 42.61% 36.78% 37.63% 47.12% 51.56% 65.52% 70.12%
GPT-3.5 0.00% 0.00% 18.73% 3.41% 4.04% 8.66% 30.82% 10.48% 38.89% 69.50% 23.71%
GPT-3.5 w/ CITYWALK 31.78% 70.24% 28.75% 12.98% 32.89% 28.72% 38.47% 27.45% 80.00% 69.94%
Table 12. Generalization Results of CITYWALK on Different LLMs in Terms of Line Coverage ($\mathbf{Cov_{L}}$)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg. $\uparrow$
CodeGeeX4 0.00% 15.30% 16.54% 0.00% 8.44% 10.45% 7.14% 21.75% 7.22% 8.11% 14.42%
CodeGeeX4 w/ CITYWALK 15.78% 46.15% 35.28% 28.04% 19.12% 12.45% 26.89% 37.89% 8.11% 9.44%
DeepSeek-V3 0.00% 0.00% 31.50% 0.00% 20.94% 22.23% 37.31% 30.04% 27.78% 16.80% 21.79%
DeepSeek-V3 w/ CITYWALK 22.44% 78.17% 42.80% 39.71% 52.23% 26.63% 48.35% 43.45% 32.87% 17.85%
GPT-3.5 0.00% 0.00% 25.59% 0.84% 10.68% 14.73% 12.09% 16.75% 22.10% 17.60% 21.67%
GPT-3.5 w/ CITYWALK 10.01% 74.87% 42.59% 30.62% 32.45% 20.12% 23.56% 32.56% 52.43% 17.85%
Table 13. Generalization Results of CITYWALK on Different LLMs in Terms of Branch Coverage ($\mathbf{Cov_{B}}$)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg. $\uparrow$
CodeGeeX4 0.00% 9.97% 15.73% 0.00% 9.18% 7.88% 10.81% 22.22% 1.35% 2.70% 10.93%
CodeGeeX4 w/ CITYWALK 5.18% 33.07% 31.23% 21.09% 21.45% 10.72% 21.53% 32.98% 5.25% 6.67%
DeepSeek-V3 0.00% 0.00% 24.24% 0.00% 19.03% 23.24% 32.58% 28.89% 15.78% 13.20% 19.30%
DeepSeek-V3 w/ CITYWALK 23.65% 65.98% 35.06% 31.94% 51.89% 25.15% 45.41% 38.12% 18.88% 13.87%
GPT-3.5 0.00% 0.00% 21.61% 0.51% 10.22% 12.50% 13.06% 13.33% 8.82% 13.32% 16.86%
GPT-3.5 w/ CITYWALK 12.70% 61.39% 34.99% 24.66% 27.12% 17.33% 19.84% 27.46% 23.09% 13.40%
4.1.3. Generalizability Evaluation
We further employ CITYWALK to three LLMs, including CodeGeeX4, DeepSeek-V3, and GPT-3.5. Specifically, we conduct ablation experiments on each LLM individually to investigate the impact of querying the corresponding LLM using the prompting strategy and additional contextual guidance designed by CITYWALK. Table 10, Table 11, Table 12, and Table 13 present the comparison results using the correctness and coverage metrics, respectively. Each LLM’s results are displayed in two lines: the first line shows the results when the LLM directly utilizes the basic prompt to generate unit test cases for the given focal methods, and the second line presents the results when integrated with CITYWALK. We derive two key insights from the statistical results: Integrating CITYWALK as a complementary plug-in enhances the performance of C++ unit test generation. As observed, the three evaluated LLMs demonstrate consistent improvements in all four metrics across the ten projects. Performance gains scale markedly with parameter count. For example, the average performance gain of each metric (listed in the Avg. $\uparrow$ column) with DeepSeek-V3-671B is substantially greater than that of CodeGeeX4-9B.
@@ 4.1.3. Generalizability Evaluation
Answer to RQ1: In conclusion, CITYWALK exhibits a marked superiority over the LLM-based baselines across all evaluation metrics, highlighting its effectiveness in C++ unit test generation. Furthermore, CITYWALK effectively harnesses the latent intelligence of the LLMs, demonstrating the potential for seamless integration with additional LLMs in a plug-and-play manner.
@@ 4.2. Answering RQ2
To answer this question, we conduct a series of ablation experiments to assess the impact of different designed components within CITYWALK. To ensure the fairness of comparisons, the parameter configurations align with those described in Section 3.5.
4.2.1. Ablation Study
@@
The components within the CITYWALK design include configuration dependencies $\mathtt{Dep_{c}}$}, cross-file data dependencies $\mathtt{Dep_{d}}$, intention contexts $\mathtt{Context_{intent}}$, guidelines of language-specific domain knowledge $\mathtt{Guideline_{DK}}$, step-by-step instructions $\mathtt{PROMPT_{step}}$, and post-processing techniques $\mathtt{Fix_{rule+prompt}}$. Specifically, we perform ablation experiments by removing one component at a time and analyze the performance contribution of each component regarding the coverage metrics $\mathbf{Cov_{L}}$ and $\mathbf{Cov_{B}}$. Table 14 and Table 15 show the performance contribution results (denoted as the performance degradation values). The greater the coverage scores decline, the larger the contribution of that component. The Avg. $\downarrow$ columns present the average results across all projects. When compared to CITYWALK, each ablation model exhibits varying degrees of decline in terms of $\mathbf{Cov_{L}}$ and $\mathbf{Cov_{B}}$, indicating that each designed component contributes to the overall enhancement in generating high-quality unit test cases. According to the statistical results presented in Table 14 and Table 15, the top three components that contribute the most to CITYWALK are $\mathtt{Dep_{c}}$}, $\mathtt{Dep_{d}}$}, and $\mathtt{Context_{intent}}$.
As discussed in Section 4.1.1, $\mathtt{Dep_{c}}$} plays a crucial role in guiding LLMs to generate syntactically correct test code by directly incorporating configuration dependencies. Thus, projects not utilizing gtest (e.g., tinyxml2 and papy) would benefit from $\mathtt{Dep_{c}}$} as it helps LLMs recognize the absence of this testing framework, thereby preventing hallucinate invocations of frameworks or libraries absent from the project’s dependencies. It is worth noting that post-processing techniques (e.g., Pynguin’s method injection (Lukasczyk and Fraser, 2022)) could also address certain types of compilation errors. In the design of CITYWALK, we integrate both proactive prevention strategies (e.g., explicitly incorporating $\mathtt{Dep_{c}}$}) and reactive post-hoc fixes (i.e., $\mathtt{Fix_{rule+prompt}}$). Our ablation study reveals that using only post-processing (w/o $\mathtt{Dep_{c}}$}) leads to a 29.49% drop in average $\mathbf{Cov_{L}}$ and a 24.58% drop in average $\mathbf{Cov_{B}}$. In contrast, using only proactive prevention (w/o $\mathtt{Fix_{rule+prompt}}$) results in a smaller decrease: 7.03% in average $\mathbf{Cov_{L}}$ and 8.60% in average $\mathbf{Cov_{B}}$. These results indicate that relying solely on post-hoc fixes is less effective than employing proactive prevention alone.
Figure 9 illustrates how extracted cross-file data dependencies are used as guidance for LLMs to generate correct test cases. By providing the key $\mathtt{Dep_{d}}$ (i.e., the initialization constructor of Json) from the json.h file, LLMs can prevent compilation errors that would otherwise arise due to invoking the non-existent function setNumber. Figure 10 illustrates how $\mathtt{Context_{intent}}$ can guide LLMs in generating correct test cases for the Failed Test Case ➋ within Figure 1. Specifically, the retrieved ParseTag method within the singledocparser class provides an invocation example of the focal method, while the retrieved _Tag method within the emittermanip header file offers an initialization example of the focal method. These examples significantly aid LLMs in understanding the usage patterns of the focal method, contributing to the generation of functionally correct assertions and enhancing code coverage.
@@
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg. $\downarrow$
CITYWALK 27.94% 77.61% 44.46% 42.56% 56.60% 25.25% 50.71% 42.15% 53.11% 24.20%
   w/o $\mathtt{Dep_{c}}$} -27.94% -77.31% -12.08% -42.56% -3.82% -1.36% -50.71% -1.81% -53.11% -24.20% -29.49%
   w/o $\mathtt{Dep_{d}}$} -0.82% -32.42% -14.50% -7.07% -15.26% -8.69% -19.04% -13.03% -15.51% -3.48% -12.98%
   w/o $\mathtt{Context_{intent}}$} -12.36% -29.84% -8.72% -8.21% -11.21% -5.11% -6.59% -9.70% -11.38% -0.90% -10.40%
   w/o $\mathtt{PROMPT_{step}}$} -5.46% -12.97% -12.33% -9.86% -8.42% -8.91% -12.48% -7.59% -16.34% -7.50% -10.19%
   w/o $\mathtt{Guideline_{DK}}$} -16.74% -7.40% -11.48% -7.87% -7.04% -2.69% -3.59% -5.59% -24.61% -6.29% -9.33%
   w/o $\mathtt{Fix_{rule+prompt}}$} -5.24% -12.51% -14.95% -8.48% -10.76% -1.51% -1.27% -4.26% -7.42% -3.94% -7.03%
@@
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg. $\downarrow$
CITYWALK 24.29% 65.66% 35.56% 35.64% 49.72% 24.64% 46.61% 37.82% 38.97% 18.05%
   w/o $\mathtt{Dep_{c}}$} -24.29% -65.66% -10.04% -35.64% -3.38% -1.50% -46.61% -1.70% -38.97% -18.05% -24.58%
   w/o $\mathtt{Dep_{d}}$} -1.99% -30.69% -11.48% -7.94% -14.60% -8.75% -19.33% -13.93% -18.61% -4.92% -13.22%
   w/o $\mathtt{Context_{intent}}$} -14.56% -28.48% -7.00% -9.87% -9.60% -5.17% -6.83% -9.59% -14.14% -2.93% -10.82%
   w/o $\mathtt{PROMPT_{step}}$} -5.97% -13.60% -10.19% -9.90% -6.16% -8.97% -12.49% -7.59% -19.04% -9.25% -10.32%
   w/o $\mathtt{Guideline_{DK}}$} -15.51% -8.07% -8.70% -9.02% -5.83% -2.75% -3.72% -5.70% -28.97% -4.65% -9.29%
   w/o $\mathtt{Fix_{rule+prompt}}$} -6.12% -9.68% -8.52% -9.60% -30.33% -3.11% -0.64% -3.26% -8.92% -5.83% -8.60%
@@
4.2.2. The Impact of Different Error-Fixing Phases within CITYWALK
We further investigate the influence of individual error-fixing phases within our three-phase post-processing techniques on test case quality enhancement. As described in Section 2.6.2, Phase ➊ and Phase ➋ focus on resolving syntactic and compilation errors via predefined rules, while Phase ➌ leverages the LLM to address compilation errors. Through ablation experiments that incrementally incorporate each phase, we quantify their respective contributions using the correctness metrics CSR and EPR. Table 16 and Table 17 present phase-by-phase performance improvements, where higher correctness score increments reflect greater phase contributions. The Avg. $\uparrow$ columns aggregate cross-project average results. Specifically, the integration of Phase ➊ yields improvements of 3.10% in CSR and 3.19% in EPR. Subsequent incorporation of Phase ➋ delivers substantial enhancements, achieving additional gains of 29.80% in CSR and 25.98% in EPR. Including Phase ➌ sustains quality improvements with further increases of 8.65% and 6.75% in CSR and EPR, respectively. The varying contributions across phases originate from differences in the types of errors resolved. Phase ➊ focuses on correcting basic syntactic errors. However, since CITYWALK already incorporates prompt optimization techniques that significantly reduce such errors, this phase yields only marginal improvements. In contrast, Phase ➋ utilizes compiler feedback to resolve a broader range of rule-based and compiler-diagnosable errors, resulting in more substantial gains. Finally, Phase ➌ targets complex compilation errors that require deeper semantic reasoning about the root causes, thereby further enhancing test case quality. Statistical analysis demonstrates that all three phases exhibit non-negative performance impacts, collectively enhancing test case quality across both metrics.
Table 16. Performance Contribution of Each Error-Fixing Phase in Terms of Compilation Success Rate (CSR)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg. $\uparrow$
No Error-Fixing Phase 77.08% 67.96% 27.59% 34.76% 22.15% 28.78% 74.56% 36.91% 54.37% 22.97%
   w/ Phase ➊ +0.00% +5.97% +0.00% +1.72% +1.60% +4.97% +0.00% +16.78% +0.00% +0.00% +3.10%
   w/ Phase ➊ ++ +22.92% +27.97% +37.32% +56.04% +44.52% +11.39% +16.58% +58.33% +0.00% +53.90% +32.90%
   w/ Phase ➊ ++++ +22.92% +29.29% +52.69% +58.81% +47.94% +19.05% +25.44% +59.79% +29.92% +69.64% +41.55%
Table 17. Performance Contribution of Each Error-Fixing Phase in Terms of Execution Pass Rate (EPR)
Project hjson-cpp tinyxml2 yaml-cpp re2 ninja leveldb json.cpp glomap papy mlx Avg. $\uparrow$
No Error-Fixing Phase 54.17% 61.27% 23.24% 31.33% 19.85% 27.12% 71.93% 27.52% 45.69% 21.70%
   w/ Phase ➊ +4.16% +7.66% +0.00% +1.58% +0.30% +3.42% +0.00% +14.76% +0.00% +0.00% +3.19%
   w/ Phase ➊ ++ +21.51% +25.61% +32.46% +51.43% +37.27% +9.44% +15.05% +47.48% +0.00% +51.43% +29.17%
   w/ Phase ➊ ++++ +23.33% +26.97% +44.58% +55.81% +42.27% +13.88% +26.98% +50.50% +7.42% +67.43% +35.92%